# Which Augmentation Should I Use? An Empirical Investigation of Augmentations for Self-Supervised Phonocardiogram Representation Learning

### v3

@@ 3.5 Implementation Details
In this section, we describe all the implementation details regarding the execution of our experiments. First of all, for the backbone encoder model we implement the same model used in ballas_listen2yourheart_2022 and papapanagiotou_chewing_2017 . The encoder model is a 5 layer convolutional neural network, the architecture of which is depicted in Figure 2 (left).
@@
For the contrastive SSL pretraining phase, we train the conv layers of the network by projecting the output of the last layer into a 128-D space via a dense or fully-connected layer and minimize the NT-Xent loss (eq. 1). For the cosine similarity (eq. 2) we use a temperature of 0.1, as in ballas_listen2yourheart_2022 , while the initial batch size is set to 256 (leading to 512 after augmentation occurs). The SSL training epochs are set to a maximum of 200 epochs, based on an early stopping set on the validation split with a patience of 10 epochs. Additionally, to accommodate the large batch size we employ the LARS optimizer ginsburg2018large with a linear learning rate warm-up of 20 epochs, up to 0.1, followed by a cosine decay of 1%.
@@
During the downstream task, the dense projection layer of the pretrained encoder is discarded and all convolutional weights are frozen. Next, a 3-layered classification head is appended to the pretrained network, as shown in Figure 2 (right). Dropout layers are also included between the dense layers of the classification head for further improving model robustness. The output size of the last decision layer depends on the downstream task type and the label size in each dataset. For the ‘Binary’ type and the PhysioNet2016 dataset, the output size is set to 1. For the ‘All’ type, the output size is set to 5 or 3, for the Pascal and PhysioNet2022 datasets respectively. In all cases, the classification head is trained via the appropriate cross-entropy loss and the Adam optimizer. Additionally, the learning rate is set to $10^{-4}$ and the batch size is set to 32. Once again, we train the model for 100 epochs at most, using early-stopping with 20 epochs patience. The same CNN model is used for the fully-supervised baseline. All experiments are implemented in Tensorflow 2.12 tensorflow2015-whitepaper and trained on a SLURM yoo_slurm_2003 cluster, containing 4$\times$40GB NVIDIA A100 GPU cards, split into 8 20GB virtual MIG devices.
4 Results and Discussion
4.1 Augmentation and Transformation Evaluation
@@
Table 1: The effect size of each augmentation on the SSL contrastive training. The effect were measured based on the out-of-distribution micro F1 scores for the PhysioNet2016 dataset, when trained on signals from PhysioNet2022, sorted by effect size. LP and HP indicate Low and High Pass filters, respectively.
@@
Cut-off (250, 200) (LP) 1.26386
Random Flip (0.3) 0.57836
Random Flip (0.7) 0.49227
Scale (1.0, 1.5) 0.20685
Noise (-0.01, 0.01) 0.11594
Inversion -0.00088
Reverse -0.00398
Scale (1.5, 2.0) -0.11695
@@
Noise (-0.001, 0.001) -0.23983
Noise (-0.1, 0.1) -0.32574
Scale (0.5, 2.0) -0.40168
Cut-off (750, 800) (HP) -0.47433
@@
Table 2: This table presents the results of the best performing models (as reported in Section 4.1)Top results are specified in bold face. The first column of the table indicates the datasets on which the SSL backbone was pretrained on. E, F, Pa, P16 and P22, indicate data from the Ephnogram, Fpcgdb, Pascal, PhysioNet16 and PhysioNet22 datasets respectively. The SSL and fully-supervised baseline models are trained and evaluated on In-Distribution data from the respective dataset. The presented models are the best-performing models, based on results from the In-Distribution test data split. Once trained, they are then evaluated on out-of-distribution data originating from unseen datasets. The ‘All’ labels indicate that the PCG signals are classified based on the dataset’s original labels (P16 has only binary labels), whereas ‘Binary’ indicates that the signals are classified between normal and abnormal. The SSL-NoDs model refers to the backbone network which has not ‘seen’ data from the downstream task dataset during pretraining.
@@
To avoid the combinatorial explosion poised by the evaluation of all possible augmentation combinations, we designed a protocol to limit the experiments to a feasible amount. As mentioned in Section 3.3, there are a total of 7 (including not transforming) possible functions that can be applied to the two copies of the audio sample, resulting in 4 possible augmentation cases; 0vs1, 1vs1, 1vs2 and 2vs2. As a first step, we examined all possible combinations in the 0vs1 and 1vs1 cases without applying the same transformation twice. After evaluating the above we selected a subset of the augmentations leading to the top performing models and continued experimenting with 1vs2 and 2vs2 augmentation schemes. The models with the best results are the ones presented in Table 2 which are discussed in the next section.
However, the main aim of this study is to provide an evaluation of the most appropriate augmentations for PCG signal analysis. Due to the fact that by only presenting the combination of the augmentations leading to the best classification results could be misleading (since only the specific combination could pertain to said results), we present a plot depicting the frequency of augmentation occurrence in the top 75 experiments across all 3 downstream tasks and for both In-Distribution and OOD data. Specifically, we select the top 25 results for each task and present the occurrences for all 75 total experiments in Fig. 3. The evaluation results are quite interesting as several conclusions can be made. First of all, the application of low-pass filters seem to be dominant in both In-Distribution and OOD top experiments. This is a clear indicator that such filters should be applied towards increasing the robustness of a classifier. What is more, the application of uniform noise, random flip and scaling on the signal, also boosts the effectiveness of classifiers in all cases. Furthermore, we found that the best performing models in each of the 3 downstream tasks, had adopted the following augmentations during their pretraining:
  • 1.
    Pascal: high pass filter (pass band 250Hz, stop band 300 Hz) and random flip (p=0.7) vs inversion and uniform noise (mean 0, std 0.1),
  • 2.
    PhysioNet2016: high pass filter (pass band 250Hz, stop band 300 Hz) vs reverse
  • 3.
    PhysioNet2022: high pass filter (pass band 250Hz, stop band 300 Hz) vs random flip (p=0.7) and random scaling with $a$ between 1.0 and 1.5
While the above augmentations agree, for the most part, with the distributions and number of occurrences in Fig. 3, they also support the findings in several previous contrastive SSL works chen_simple_2020 ; gopal_3kg_2021 , in which stronger augmentations lead to improved results.
To further validate our results and to draw concrete conclusions regarding the importance of each applied transformation, we calculate Cohen’s $d$ cohen2013statistical for each implemented function. Cohen’s $d$ provides a standardized measure of the effect size of each augmentation and allows the comparison of their magnitude. Given two groups $x_{1}$ and $x_{2}$, Cohen’s $d$ is the difference in their means divided by the pooled standard deviation ($s$). Specifically, Cohen’s $d$ is defined as:
$d=\frac{\bar{x_{1}}-\bar{x_{2}}}{s}$ (3)
where the pooled standard deviation is defined as:
$s=\sqrt{\frac{(n_{1}-1){s^{2}}_{1}+(n_{2}-1){s^{2}}_{2}}{n_{1}+n_{2}-2}}$ (4)
and $n_{1}$, $n_{2}$ are the variances for each group respectively. The consensus on the interpretation of Cohen’s d is generally that its sign means to either a negative or positive effect, while a magnitude of $d=0.01$ up to $d=2.0$ corresponds to a very small or very large effect, respectively. To calculate Cohen’s $d$, we create two subgroups for each augmentation. Specifically, let $a$ be an augmentation under consideration. The first group ($x_{1}$) contains all experiments where $a$ was implemented, while the second group ($x_{2}$) consists of all experiments with the same set of augmentations but without $a$. For example, say that the augmentation under consideration is Noise. One experiment to include in the first group would be the Noise vs Inversion experiment. Therefore the corresponding experiment which should be added to the second group would be None vs Inversion. The same process can be repeated for all augmentations and for all runs, including $22vs$ 2 experiments, as specified in Sec. 3.3. Accordingly, we calculate the effect of each augmentation based on the Out-of-Distribution Micro F1 scores for the PhysioNet2016 dataset, when trained on signals from PhysioNet2022. The magnitude of effect regarding each augmentation is presented in Table 1.
Interestingly enough, the statistics presented in Fig. 3, seem to agree with the effect sizes of each augmentation (Table 1). Based on Cohen’s $d$ measure, the low-pass filters have the most important positive effect on robust SSL PCG learning. In simple terms, this means that the addition or application of low-pass cut-off filters during the pretraining phase of the encoder plays a significant role ($d>1.0$) in learning robust representations. Additionally, the application of Reverse and Inversion (i.e Random Flip) has an overall positive effect on training, as shown by the positive $d$ measure. What’s more, the balanced reversion of inversion of the signals (i.e Random Flip with a .5% chance), has the largest positive effect ($d\approx 1.8$), whereas the application of only one of the above flip augmentations seems to have approximately no effect on training. Similarly, but to a less degree, scaling the signals by a small factor, assists the model in extracting meaningful representations. On the other hand, the addition of noise, the application of high-pass filters and large scaling, generally appear to have a negative effect. The intuition behind the negative effect of noise, is that since the selected signals derive from an already “noisy” dataset, the additive noise renders them meaningless.
In the next section we present numerical results for all downstream tasks and experiments.
4.2 Results
@@
Table 2 presents the Accuracy and F1-Scores of the best performing models trained via contrastive SSL for all downstream tasks, including evaluation on OOD distribution data, as presented in Section 3.4. We perform a total of 4 cycles of SSL experiments, evaluating all augmentation combinations in each cycle. In each experiment cycle, we either use all available datasets or leave one labeled dataset out at a time. The first column of Table 2 denotes the datasets used during SSL training. In the same table, the ‘Baseline’ model refers to the fully-supervised model trained solely on data from the downstream ‘In-Distribution’ dataset. Due to the variance in the OOD results for the baseline model, we present the average of 5 total runs. Finally, the ‘SSL-NoDs’ model refers to the model which has not been pretrained with data originating from the dataset of the downstream task. All results are rated based on the F1-Scores on the ‘In-Distribution’ test splits and the adopted augmentations for each case are the ones mentioned in Section 4.1. The same models are then evaluated on the test splits of the OOD datasets.
4.3 Discussion
@@
By diving further into the results, the true benefits of the proposed method begin to show. First of all, the drop in effectiveness between in-distribution and OOD data is apparent. Additionally, in all cases the baseline models appear to have overfit their training data distribution and fail to maintain their classification ability across distinct datasets. On the other hand, the models which have been pretrained via SSL yield far superior results. In all cases, the models trained with the proposed framework surpass the baseline by 11.68% on average and even outperform it by 21.04% in the case of the PhysioNet22 OOD task. Even more so, the SSL models were surprisingly able to surpass both the effectiveness of the baseline and their ‘In-Distribution’ counterpart, in the case of the Pascal OOD evaluation. Specifically, the downstream models trained on data from PhysioNet2016 and PhysioNet2022 were able to yield better results than the fully-supervised model trained solely on Pascal data. Furthermore, even though there is a decrease in effectiveness when the OOD datasets are left out during model pretraining, the SSL models continue to outshine the baseline across the board, with the exception of the case where the downstream task is the PhysioNet2022 data and models are evaluated on OOD data from Pascal. Finally, we would like to note that the above findings are aligned with the results of our preliminary SSL method ballas_listen2yourheart_2022 submitted in the PhysioNet2022 challenge, where the proposed model was able to avoid overfitting on the available training distribution and demonstrated adequate generalization effectiveness on completely hidden data.
5 Conclusions
This research presents a comprehensive and extensive evaluation of augmentation and transformations for pretraining neural networks, towards robust PCG classification. In addition to presenting the best performing combinations of augmentations, we also provide a summary of the most frequently appearing transformation functions in the top 75 overall experimental results. Our experiments suggest that the use of cutoff filters, particularly low pass filters, leads to the best performing models among several datasets, rendering them almost essential during such tasks. Furthermore, the addition of small random noise and inversion or reversal of the signal, also yields improved results. Finally, the numerical results of our experiments validate our claims that contrastive SSL learning can induce robust and generalizable classifiers, as the drop in effectiveness when evaluating on previously unseen data is far less than that of a fully-supervised model. The proposed method however, does not come without drawbacks. Most importantly, to pretrain a model via SSL, most likely a large corpus of data must be available beforehand in order to enhance the representation extraction capabilities of the model. What’s more, the implemented contrastive loss expects a large batch size to perform as expected, adding a computational constraint to the overall framework. In the future, we aim to enhance the adopted augmentation techniques, by adding novel functions or researching further parameters of the existing ones, expanding the downstream tasks by incorporating transfer learning problems, as well as applying it to other 1D biosignals such as EMG.
Acknowledgement
The work leading to these results has been funded by the European Union under Grant Agreement No. 101057821, project RELEVIUM.

### v4

@@ III-E Implementation Details
In this section, we describe all the implementation details regarding the execution of our experiments. First of all, for the backbone encoder model we implement the same model used in [27] and [72]. The encoder model is a 5 layer convolutional neural network, the architecture of which is depicted in Figure 2 (left). Given a batch size of $b$ the input of the encoder model is a tensor of dimension $(b,10000)$.
@@
For the contrastive SSL pretraining phase, we train the conv layers of the network by projecting the flattened output of the last layer, with an original dimension of $(b,8,64)$ into a 128-D space via a dense or fully-connected layer and by ultimately minimizing the NT-Xent loss (eq. 1). The dimension of the projection was set to follow the one in the original SimCLR paper [70]. For the cosine similarity (eq. 2) we use a temperature of 0.1, as in [27], while the initial batch size is set to 256 (leading to 512 after augmentation occurs). The SSL training epochs are set to a maximum of 200 epochs, based on an early stopping set on the validation split with a patience of 10 epochs. Additionally, to accommodate the large batch size we employ the LARS optimizer [73] with a linear learning rate warm-up of 20 epochs, up to 0.1, followed by a cosine decay of 1%.
@@
During the downstream task, the dense projection layer of the pretrained encoder is discarded and all convolutional weights are frozen. Followingly, a 3-layered classification head is appended to the pretrained network, as shown in Figure 2 (right). The dimensions of each layer are set to reflect the model implemented in [27], which was shown to yield promising results. Dropout layers are included between each of the dense layers of the classification head in order to avoid overfitting. The application of dropout between multiple fully connected layers has shown improve the generalizability and robustness of classifiers [74, 72]. The output size of the last decision layer depends on the downstream task type and the label size in each dataset. For the ‘Binary’ type and the PhysioNet2016 dataset, the output size is set to 1. For the ‘All’ type, the output size is set to 5 or 3, for the Pascal and PhysioNet2022 datasets respectively. In all cases, the classification head is trained via the appropriate cross-entropy loss and the Adam optimizer. Additionally, the learning rate is set to $10^{-4}$ and the batch size is set to 32. Once again, we train the model for 100 epochs at most, using early-stopping with 20 epochs patience. The same CNN model is used for the fully-supervised baseline. All experiments are implemented in Tensorflow 2.12 [75] and trained on a SLURM [76] cluster, containing 4$\times$40GB NVIDIA A100 GPU cards, split into 8 20GB virtual MIG devices.
IV Results and Discussion
IV-A Augmentation and Transformation Evaluation
@@
TABLE I: The effect size of each augmentation on the SSL contrastive training. The effect were measured based on the out-of-distribution micro F1 scores for the PhysioNet2016 dataset, when trained on signals from PhysioNet2022, sorted by effect size. LP and HP indicate Low and High Pass filters, respectively.
@@
Cut-off (250, 200) (LP) 1.2639
Random Flip (0.3) 0.5784
Random Flip (0.7) 0.4923
Scale (1.0, 1.5) 0.2068
Noise (-0.01, 0.01) 0.1159
Inversion -0.0008
Reverse -0.0039
Scale (1.5, 2.0) -0.1169
@@
Noise (-0.001, 0.001) -0.2398
Noise (-0.1, 0.1) -0.3257
Scale (0.5, 2.0) -0.4016
Cut-off (750, 800) (HP) -0.4743
To avoid the combinatorial explosion poised by the evaluation of all possible augmentation combinations, we designed a protocol to limit the experiments to a feasible amount. As mentioned in Section III-C, there are a total of 7 (including not transforming) possible functions that can be applied to the two copies of the audio sample, resulting in 4 possible augmentation cases; 0vs1, 1vs1, 1vs2 and 2vs2. As a first step, we examined all possible combinations in the 0vs1 and 1vs1 cases without applying the same transformation twice. After evaluating the above we selected a subset of augmentations which led to the top performing models in the “In-Distribution” setting and continued experimenting with 1vs2 and 2vs2 augmentation schemes. The models with the best results are presented in Table II and are discussed in the next section.
The main aim of this study is to provide an evaluation of the most appropriate augmentations for PCG signal analysis. Due to the fact that by only presenting the combination of the augmentations leading to the best classification results could be misleading (since only the specific combination could pertain to said results), we present a plot depicting the frequency of augmentation occurrence in the top 75 experiments across all 3 downstream tasks and for both In-Distribution and OOD data. Specifically, we select the top 25 results for each task and present the occurrences for all 75 total experiments in Fig. 3. The evaluation results are quite interesting as several conclusions can be drawn. First of all, the application of low-pass filters seem to be dominant in both In-Distribution and OOD top experiments. This is a clear indicator that such filters should be applied towards increasing the robustness of a classifier. What is more, the application of uniform noise, random flip and scaling on the signal, also boosts the effectiveness of classifiers in all cases. Furthermore, we found that the best performing models in each of the 3 downstream tasks, had adopted the following augmentations during their pretraining:
  • Pascal: low pass filter (pass band 250Hz, stop band 300 Hz) and random flip (p=0.7) vs inversion and uniform noise (mean 0, std 0.1),
  • PhysioNet2016: low pass filter (pass band 250Hz, stop band 300 Hz) vs reverse
  • PhysioNet2022: low pass filter (pass band 250Hz, stop band 300 Hz) vs random flip (p=0.7) and random scaling with $a$ between 1.0 and 1.5
While the above augmentations agree, for the most part, with the distributions and number of occurrences in Fig. 3, they also support the findings in several previous contrastive SSL works [66, 77], in which stronger augmentations lead to improved results.
To further validate our results and to draw concrete conclusions regarding the importance of each applied transformation, we calculate Cohen’s $d$ [71] for each implemented function. Cohen’s $d$ provides a standardized measure of the effect size of each augmentation and allows the comparison of their magnitude. Given two groups $x_{1}$ and $x_{2}$, Cohen’s $d$ is the difference in their means divided by the pooled standard deviation ($s$). Specifically, Cohen’s $d$ is defined as:
$d=\frac{\bar{x_{1}}-\bar{x_{2}}}{s}$ (3)
where the pooled standard deviation is defined as:
$s=\sqrt{\frac{(n_{1}-1){s^{2}}_{1}+(n_{2}-1){s^{2}}_{2}}{n_{1}+n_{2}-2}}$ (4)
and $n_{1}$, $n_{2}$ are the variances for each group respectively. The consensus on the interpretation of Cohen’s d is generally that its sign means to either a negative or positive effect, while a magnitude of $d=0.01$ up to $d=2.0$ corresponds to a very small or very large effect, respectively. To calculate Cohen’s $d$, we create two subgroups for each augmentation. Specifically, let $a$ be an augmentation under consideration. The first group ($x_{1}$) contains all experiments where $a$ was implemented, while the second group ($x_{2}$) consists of all experiments with the same set of augmentations but without $a$. For example, say that the augmentation under consideration is Noise. One experiment to include in the first group would be the Noise vs Inversion experiment. Therefore the corresponding experiment which should be added to the second group would be None vs Inversion. The same process can be repeated for all augmentations and for all runs, including $22vs$ 2 experiments, as specified in Sec. III-C. Accordingly, we calculate the effect of each augmentation based on the Out-of-Distribution Micro F1 scores for the PhysioNet2016 dataset, when trained on signals from PhysioNet2022. The magnitude of effect regarding each augmentation is presented in Table I.
Interestingly enough, the statistics presented in Fig. 3, seem to agree with the effect sizes of each augmentation (Table I). Based on Cohen’s $d$ measure, the low-pass filters have the most important positive effect on robust SSL PCG learning. In simple terms, this means that the addition or application of low-pass cut-off filters during the pretraining phase of the encoder plays a significant role ($d>1.0$) in learning robust representations. The above result agrees with the general consensus that important biosignal artifacts are present in low frequencies, whearas unrelated noise is attributed to higher frequencies [78, 79, 80]. Additionally, the application of Reverse and Inversion (i.e Random Flip) has an overall positive effect on training, as shown by the positive $d$ measure. What’s more, the balanced reversion of inversion of the signals (i.e Random Flip with a .5% chance), has the largest positive effect ($d\approx 1.8$), whereas the application of only one of the above flip augmentations seems to have approximately no effect on training. Similarly, but to a less degree, scaling the signals by a small factor, assists the model in extracting meaningful representations. On the other hand, the addition of noise, the application of high-pass filters and large scaling, generally appear to have a negative effect. The intuition behind the negative effect of noise, is that since the selected signals derive from an already “noisy” dataset, the additive noise renders them meaningless, while the application of high-pass filters removes important signal artifacts present in lower frequencies.
Even though low-pass filters seem to be have the largest postive effect in training, they have to be combined with further transformations for a SSL model to reach optimal performance, as already mentioned. In the next section we present numerical results for all downstream tasks and experiments.
(a) P16 - Baseline
(b) P16 - SSL
(c) P22 (OOD) - Baseline
(d) P22 (OOD) - SSL
Figure 4: T-SNE visualizations of feature vectors extracted from the penultimate layer, after model training of a baseline fully-supervised model and a model trained via contrastive SSL. In order to exhibit the generalization capabilities of a model trained via the proposed framework, we plot the features of both In-Distribution, from the PhysioNet2016 dataset and OOD samples from the PhysioNet2022. Abnormal signals are colored in blue, while normal signals are shown in orange. As illustrated, the features extracted by the baseline models are chaotic and perplexed, whereas the SSL model seems to group samples in a more structured manner.
IV-B Results
@@
TABLE II: This table presents the results of the best performing models (as reported in Section IV-A)Top results are specified in bold face. The first column of the table indicates the datasets on which the SSL backbone was pretrained on. E, F, Pa, P16 and P22, indicate data from the Ephnogram, Fpcgdb, Pascal, PhysioNet16 and PhysioNet22 datasets respectively. The SSL and fully-supervised baseline models are trained and evaluated on In-Distribution data from the respective dataset. The presented models are the best-performing models, based on results from the In-Distribution test data split. Once trained, they are then evaluated on out-of-distribution data originating from unseen datasets. The ‘All’ labels indicate that the PCG signals are classified based on the dataset’s original labels (P16 has only binary labels), whereas ‘Binary’ indicates that the signals are classified between normal and abnormal. The SSL-NoDs model refers to the backbone network which has not ‘seen’ data from the downstream task dataset during pretraining.
@@
Table II presents the Accuracy and F1-Scores of the best performing models trained via contrastive SSL for all downstream tasks, including evaluation on OOD distribution data, as presented in Section III-D. We perform a total of 4 cycles of SSL experiments, evaluating all augmentation combinations in each cycle. In each experiment cycle, we either use all available datasets or leave one labeled dataset out at a time. The first column of Table II denotes the datasets used during SSL training. In the same table, the ‘Baseline’ model refers to the fully-supervised model trained solely on data from the downstream ‘In-Distribution’ dataset. Due to the variance in the OOD results for the baseline model, we present the average of 5 total runs. Finally, the ‘SSL-NoDs’ model refers to the model which has not been pretrained with data originating from the dataset of the downstream task. All results are rated based on the F1-Scores on the ‘In-Distribution’ test splits and the adopted augmentations for each case are the ones mentioned in Section IV-A. The same models are then evaluated on the test splits of the OOD datasets.
IV-C Discussion
@@
By diving further into the results, the true benefits of the proposed method begin to show. First of all, the drop in effectiveness between in-distribution and OOD data is apparent. Additionally, in all cases the baseline models appear to have overfit their training data distribution and fail to maintain their classification ability across distinct datasets. On the other hand, the models which have been pretrained via SSL yield far superior results. In all cases, the models trained with the proposed framework surpass the baseline by 11.68% on average and even outperform it by 21.04% in the case of the PhysioNet22 OOD task. Even more so, the SSL models were surprisingly able to surpass both the effectiveness of the baseline and their ‘In-Distribution’ counterpart, in the case of the Pascal OOD evaluation. Specifically, the downstream models trained on data from PhysioNet2016 and PhysioNet2022 were able to yield better results than the fully-supervised model trained solely on Pascal data. Furthermore, even though there is a decrease in effectiveness when the OOD datasets are left out during model pretraining, the SSL models continue to outshine the baseline across the board, with the exception of the case where the downstream task is the PhysioNet2022 data and models are evaluated on OOD data from Pascal. Finally, we would like to note that the above findings are aligned with the results of our preliminary SSL method [27] submitted in the PhysioNet2022 challenge, where the proposed model was able to avoid overfitting on the available training distribution and demonstrated adequate generalization effectiveness on completely hidden data.
To further demonstrate the generalization capabilities of the SSL models, we provide t-SNE visualizations and illustrate the feature distribution of representations extracted from the penultimate layer of the best-performing model, in addition to representations extracted from a fully-supervised baseline model. Specifically, we visualize the features from signals drawn from distributions the models have been trained on (In-Distribution) and unseen distributions (OOD) in Fig. 4. As depicted, the feature distribution of the baseline models are disordered, without a clear distinction between the abnormal (blue) and normal (signals). On the other hand, the features of the SSL model portray a somewhat clearer structure with the features of each signal type grouping into clusters towards the upper-right and bottom-left corners. The compelling aspect of the visualization results, lies in the fact that the general feature distributions of the SSL models align on both in-distribution and completely unseen during training OOD samples.
V Conclusions & Future Work
This work provides a thorough evaluation of augmentation techniques and transformations for pretraining neural networks to achieve robust phonocardiogram (PCG) classification. Our method focuses on identifying the most effective augmentations for improving the performance of self-supervised learning (SSL) models. We present both the top-performing combinations of augmentations and a summary of the most frequently used transformation functions in the top 75 experimental results. Our findings highlight that the use of low-pass cutoff filters, consistently leads to the best-performing models across various datasets, making them essential for this task. Additionally, adding small random noise and applying signal inversion or reversal further enhance model performance.
Our results validate that contrastive SSL significantly improves the robustness and generalizability of PCG classifiers. Compared to fully-supervised models, SSL models experience a much smaller drop in effectiveness when evaluated on unseen data. However, the proposed method has limitations: SSL pretraining requires access to a large data corpus to maximize the model’s representation learning capabilities, and the contrastive loss function relies on large batch sizes, imposing a computational constraint.
In future work, we plan to refine the augmentation techniques by introducing new functions and exploring additional parameters, extend our approach to transfer learning tasks, and investigate the most effective augmentation policies for other 1D biosignals such as ECG, EMG, and EEG classification.